OpenStudy (anonymous):

1. If C = {integers divisible by 2 from 1 to 12} and D = {integers divisible by 4 from 1 to 16}, what is C ∩ D? A{4, 8, 12} B{2, 4, 6, 8, 10, 12, 16} C{8, 12} D{ } 2. The range of the relation below is x y −1 10 2 10 5 4 0 3 A {3, 4, 10} B {−1, 0, 2, 5} C {10, 10, 4, 3} D No range exists. 3.Given the function f(x) = 2(x + 6), find x if f(x) = 22. A 56 B 5 C 38 D 17

OpenStudy (ganpat):

1. C ∩ D = all the common elements from C & D... 2. confusing.. 3. F(x) = 22 = 2 (x +6)... can u find for x ?? can u give a try ??

OpenStudy (ganpat):

my mother ?? can try or not ??

OpenStudy (ganpat):

If C = {integers divisible by 2 from 1 to 12} = {2,4,6,8,10,12} and D = {integers divisible by 4 from 1 to 16} = {4,8,12,16} what is C ∩ D = all common elements in C & D = {4,8,12} so option A

OpenStudy (ganpat):

3. f(x) = 2(x + 6), f(x) = 22. so f(x) = 2(x + 6)= 22. 2x +12 = 22 2x = 10 x = 5 option B

OpenStudy (ganpat):

for B, i dont find any range.. question is not very clear for me..